there are 4 types of toys (a,b,c,d) . There is one toy in each cereal box. A mother buys 6 boxes. Determine the probability that she is able to collect all 4 toys
1
Expert's answer
2012-09-25T11:34:56-0400
Total possible combinations of distribution toys in boxes is
4^6 = 4096
Number of successful combinations(all types of toys)
4 *4 *4 *3 *2 *1 = 384
The probability is
384/4096 = 3/32 = 0,09375